OpenStudy (pawanyadav):

How to integrate cos(t^2) from 0 to 1

OpenStudy (nuttyliaczar):

First set up the integral with bounds from 0 to 1 with infinitesimal dt. Then just take the derivative in reverse (what gives you cos(t^2) when you take the derivative of it?)

zepdrix (zepdrix):

Hey Pawanyadav :) This does not have a closed form solution. Are you sure you posted the question correctly? Or are you using approximation methods maybe? Or power series or something? :d

OpenStudy (nuttyliaczar):

Oh ew you're right, this problem has cooties

zepdrix (zepdrix):

lol :3

OpenStudy (pawanyadav):

That is Cosx/√x dx So by putting. x=t^2 I get this

OpenStudy (pawanyadav):

@zepdrix,@nuttyliaczar

zepdrix (zepdrix):

Oh ok that makes more sense :) I think maybe you missed something in your substitution.

zepdrix (zepdrix):

\[\large\rm \int\limits\frac{\cos(x)}{\sqrt x}dx=\int\limits \cos(x)\left(\frac{1}{\sqrt x}dx\right)\]

zepdrix (zepdrix):

\[\large\rm x=t^2\qquad\to\qquad \sqrt{x}=t\]Taking derivative,\[\large\rm dx=2t~dt\qquad\to\qquad dx=2\sqrt x~dt\]

OpenStudy (nuttyliaczar):

So to be clear the function is cos(x/sqrt(x))? With x=t^2

zepdrix (zepdrix):

So then you have:\[\large\rm \left(\frac{1}{\sqrt x}dx\right)=2 dt\]Oh no, I guess you did that correctly :)\[\large\rm \int\limits\limits \cos(x)\left(\frac{1}{\sqrt x}dx\right)=\int\limits \cos(t^2)\left(2dt\right)\]Mmm that's no fun >.<

OpenStudy (pawanyadav):

I take 2 outside the integral

zepdrix (zepdrix):

Hmm yah I dunno how to deal with this one using normal methods :( Here is what wolfram provides https://www.wolframalpha.com/input/?i=integral+of+cos%28x%29%2Fsqrtx you have to use some other weird trick I think.

OpenStudy (pawanyadav):

Am not getting anything